Hello, Im stuck can i please have some help. Thank you, kindly!
In 1995 the Life Expectancy of a certain country was 68.9 years. In 2002 it was 71.1 years. Let E represent the life expectancy in year T and let year T represent the number of years since 1995.
E(t) = __t+___
(round to the nearest tenth)
Use the function to predict the life expectancy in males in 2005.
E(10)= _____
(Round to nearest tenth)

Answer by nerdybill(7384)

In 1995 the Life Expectancy of a certain country was 68.9 years. In 2002 it was 71.1 years. Let E represent the life expectancy in year T and let year T represent the number of years since 1995.
E(t) = __t+___
(round to the nearest tenth)
This is really in "slope-intercept" form.
The problem gives you two points (assume 1995 is zero):
(0, 68.9)
(7, 71.1)
slope = (y2 - y1)/(x2 - x1)
slope = (71.1 - 68.9)/(7 - 0)
slope = 2.2/7
slope = .3
.
Now, using one point say (0, 68.9) and the slope (.3) plug it into the "point-slope" form:
y-y1 = m(x-x1)
y-68.9 = .3(x-0)
y-68.9 = .3x
y = .3x+68.9
.
So,
E(t) = __t+___
becomes
E(t) = .3t+68.9
.
Use the function to predict the life expectancy in males in 2005.
E(10)= _____
(Round to nearest tenth)
E(t) = .3t+68.9
E(10) = .3(10)+68.9
E(10) = 3+68.9
E(10) = 71.9 years